5. Using regression analysis to forecast assets

The AFN equation and the financial statement–forecasting approach both assume that assets grow at relatively the same rate as sales. However, the relationship between assets and sales is often a little more difficult than that. In particular, some firms use regression analysis to predict the required assets needed to support a given level of sales.

Mile Brewing Co. has used its historical sales and asset data to estimate the following regression equations:

Accounts Receivable = –$95,420 + 0.249(Sales)
Inventories = $8,765 + 0.243(Sales)

Mile Brewing Co. currently has sales of $922,000, but it expects sales to grow by 20% over the next year. Use the regression models to calculate Mile Brewing Co.’s forecasted values for accounts receivable and inventories needed to support next year’s sales.

Forecasted Values for Next Year

Accounts receivable??????????
Inventories????????????

Based on the next year’s accounts receivable and inventory levels predicted by Mile Brewing Co.’s regression equations, the firm’s DSO for next year is expected to be??????????    . Use 365 days as the length of a year in all calculations.
Answer #1

Forecasted value

Accounts receivable = -95,420 + 0.249*(922000*120%)

= 370913.60

Inventories= 8765+0.243*(922000*120%)

=277620.20

Days sales outstanding = Receivables*365/Sales

= 370913.60*365/(922000*120%)

=122.36